A small square is located in side a big square The lentgh of one side of the small square is 3 inches and the length of the big square is 7 inches What is the area of he region located outside the small square but inside the big square?
Answer by macston(5194)

.
Area of large square=(7in)^2=49in^2
Area of small square=(3in)^2=9in^2
Large square - small square=49in^2-9in^2=40in^2
ANSWER: The area inside the large square outside the small square is 40 square inches.